# LLM-Based Program Generation for Triggering Numerical Inconsistencies Across Compilers

### v1

@@ 6. Conclusion
We propose LLM4FP, a novel LLM-based approach for generating floating-point programs that trigger numerical inconsistencies across compilers and/or optimization levels. By combining Grammar-Based Generation and Feedback-Based Mutation, LLM4FP significantly outperforms existing tools like Varity, detecting over twice as many inconsistencies. Our evaluation shows improved program diversity over other LLM-based baselines with comparable time cost. Most inconsistencies it finds involve real-valued results, and they demonstrate LLM4FP’s effectiveness at exposing inconsistencies across a wide range of optimization levels and between host and device compilers. In future work, we plan to enhance semantic guidance and integrate static analysis to improve program quality and better distinguish harmful inconsistencies from benign ones.

### v2

@@ 6. Conclusion
We propose LLM4FP, a novel LLM-based approach for generating floating-point programs that trigger numerical inconsistencies across compilers and/or optimization levels. By combining Grammar-Based Generation and Feedback-Based Mutation, LLM4FP significantly outperforms existing tools like Varity, detecting nearly 2.5 times as many inconsistencies. Our evaluation shows improved program diversity over other baselines, with a trade-off in time cost. Most inconsistencies it finds involve real-valued results, and they demonstrate LLM4FP’s effectiveness at exposing inconsistencies across a wide range of optimization levels and between host and device compilers. Our code, documentation and experimental data are publicly available at https://github.com/ucd-plse/LLM4FP.